# An interpretable generative multimodal neuroimaging-genomics framework for decoding Alzheimer’s disease

### v1

@@ 4.3 Group-based statistical IG analysis
The IG scores obtained for the four groups of subjects in the study were analyzed together in order to quantify the differences and similarities outlined in the qualitative analysis. In what follows, the statistical analyses will be presented separately for the neuroimaging (sMRI and rs-fMRI) and the genetic channels.
Neuroimaging channels
@@
Figure 7: Enrichment analysis results. The top $15$ biological processes are reported for the three classes of patients, namely AD, MCIc, and MCInc (columns). The enrichment analysis was performed over the genes annotated from the SNPs-IG. The colorbar represents the associated p-value, dark red the lowest, while the number of genes involved is reported as the bar height. The * symbol indicates the biological processes statistically significant after p-value correction.
@@
Fig. 6A shows the boxplots representing the distribution of the average IG values for each subject in the most relevant cortical and subcortical brain regions (ROIs). Of note, only those regions where at least the $99\%$ of the subjects had an IG value were selected for these analyses. The complete acronym list is reported in Supplementary Table 1. A general agreement between the CN and the MCInc subjects, as well as between the AD and the MCIc patients is evident in all the considered ROIs, with the AD/MCIc subjects showing a generally higher variance. As expected from the qualitative analysis, the Hipp and the Amy resulted as the highest relevant ROIs, as well as the ones showing the highest distance between the groups, with high positive attribution for the CN and the MCInc, and strong negative attribution for the AD and MCIc. Among the other subcortical ROIs, the Acc showed a notable relevance for all the groups, with the MCInc and CN having negative scores and the MCIc and AD positive ones. Cau and Put had an associated high positive relevance for AD and MCIc, while almost null scores for the other two groups. Interestingly, among the cortical regions, the TOF was the most relevant region for both the MCIc and AD (positive IG scores) and the MCInc and CN (negative IG scores). The highest distance between the group IG scores distributions was found for PhGa, followed by the TFCp with the AD and MCIc having high negative scores. On the contrary, the OFG, the POpC, and the PaGp showed high positive relevance for AD and MCIc patients associated with negative scores for the CN and MCInc.
In order to evaluate whether these differences were significant, a Kruskal-Wallis test was performed separately for each ROI, considering the four groups together. The significant ROIs (all except the Ins and the TP) were further analyzed to investigate the group-related differences through the Wilcoxon Rank Sum test. Results are shown in Fig. 6B. The heatmap reports the Bonferroni corrected p-values ($6$ pairwise comparisons) in logarithmic scale and reverse hot colorbar (yellow/red correspond to higher/lower p-value, white stands for the non-significant comparisons). The most significant differences were found between MCIc and MCInc in the subcortical ROIs, resulting as generally significant (except for Put and Pall) with the lowest p-values being recorded for Hipp and Amy. Moreover, significant differences were also recorded for the temporal (PcC, PhGa, PaGp, TFCp) and occipital (TOF, OFG, POpC) cortical regions, with the most significant difference being recorded for PcC and PhGa. A coherent significance pattern was found between the contrasts MCIc-CN, MCInc-AD, and the CN-AD with generally higher p-values. Of interest, no statistically significant differences were found in the sMRI-IG scores distributions between AD and MCIc patients, except for PcC, TFCp, and POpC, while some significant differences were found between CN and MCInc but with relatively high p-values.
@@
The same statistical analysis was carried out for fMRI-IG, considering all the most relevant connections resulting from the fMRI-IG scores for all the considered classes. The connections between IC18181818 (VI) – IC14141414 (SM) and IC44444444 (DM) – IC17171717 (VI) did not reach significance at the Kruskal-Wallis test, therefore they were excluded from the post-hoc analysis. The Bonferroni-corrected pairwise tests (Wilcoxon rank sum) are reported in Fig. 6C, separately for intra- and inter-network connections, with the p-values being represented in logarithmic scale. The related boxplots are reported in Supplementary Fig. $1$. Opposed to the sMRI-IG significance pattern, the most significant differences were found between the CN-AD and MCInc-AD contrasts, which showed significant differences for almost all the connections considered. In particular, for the MCInc-AD contrast, significant intra-network connections were part of the SM, involving connections in the parietal lobe, and in the VI RSNs. Concerning the inter-network connection, significance was found between ICs in the DM, part of the cingulate gyrus (IC47, IC45), and ICs in the CC or VI RSNs, as well as between ICs in the SM, located in the parietal lobe and ICs in the CC or CB RSNs. MCIc-CN and MCInc-MCIc showed the same significant pattern as the CN-AD contrast, showing generally higher p-values. Of interest, as for the sMRI-IG statistical analysis, no significant differences were detected for the contrasts MCIc-AD and MCInc-CN, except for two connections between CB and SM (IC51-IC9 and the IC51-IC12 respectively for the MCInc-CN and the MCIc-AD).
Genetic channel
@@ 5 Discussion
With this work, we proposed a novel multimodal generative and interpretable method, which holds the potential of addressing missing data management while focusing on model interpretability. We specifically applied this method to the crucial tasks of diagnosing AD and assessed its viability in detecting also MCI conversion to AD. By integrating neuroimaging (sMRI and rs-fMRI) and genetics (SNPs) data through DL-based feature extraction, we introduced a novel multimodal data fusion framework, relying on input data not yet simultaneously investigated in the AD continuum in the current literature. Incomplete data is handled by generating missing modalities in the latent space, obtained after feature reduction. This ensures high accuracy in reconstructing latent features while minimizing computational demands. To accomplish this, we relied on pre-trained generators from two cGANs, one trained to generate sMRI and SNPs data, and the other to generate sMRI and rs-fMRI modalities. The approach is however generalizable to any missing modality. The framework is trained to detect AD in a cohort including CN and AD subjects, while its generalization capabilities were then tested by predicting MCI conversion to AD in an independent study cohort. Furthermore, we meticulously conducted a post-hoc interpretability analysis, supplemented by a robust validation step, consolidating the findings’ impact in the field. In this section we will first give a technical discussion focused on the classification accuracies and the missing data management, we will then discuss the interpretability analysis and the obtained results from a more neuroscientific point of view.
Table 3: Comparison of the proposed model with other state-of-the-art methods dealing with missing modalities for AD detection and MCI conversion tasks.
(a) AD detection task. Accuracy (ACC), precision (PRE), and recall (REC) metrics are reported on the test set or averaged during the cross-validation phase ($mean\pm std$).
Authors Modalities Study cohort Input data Missing data rate Missing data management ACC REC PRE XAI [33] sMRI, Genetics, Clinics 1406 PAT (266 AD, 699 MCI), 598 CN Full MRI volume WGS data, Set of clinical scores $\sim 67\%$ missing sMRI $\sim 61\%$ missing SNP Feature exclusion if the 70% had missing value Zero filling for the others $0.630^{*}$ $0.780^{\dagger}$ $0.570^{*}$ $0.780^{\dagger}$ $0.620^{*}$ $0.770^{\dagger}$ Yes [35] sMRI, FDG-PET 85 AD, 90 CN ROI based features $50\%$ missing PET Auto-Encoder for the complementation of complete and incomplete latent representation $0.836\pm 0.08$ n.d. n.d. No [38] sMRI, FDG-PET 352 AD, 427 CN sMRI and PET volumes $30\%$ missing PET Task-induced pyramid and attention generative adversarial network (TPA-GAN) $0.927$ $0.917$ n.d. No [64] sMRI, FDG-PET 160 AD, 210 CN ROI based features for VBM, FDG 50% missing PET GAN with attention layer in latent space $0.914\pm 0.19$ $0.881\pm 0.11$ n.d. Yes Proposed framework sMRI, fMRI, Genetics $332$ AD, $664$ CN GM volume sFNC matrix, SNPs $\sim 59\%$ missing fMRI $\sim 57\%$ missing SNPs Multiple cGANs to generate missing latent representation $0.926\pm 0.02$ $0.876\pm 0.03$ $0.910\pm 0.05$ Yes
(b) MCI conversion prediction. Accuracy (ACC), precision (PRE), and recall (REC) metrics are reported on the test set.
Authors Modalities Study cohort Input data Missing data rate Missing data management Trained on the specific task ACC REC PRE XAI [34] $10$ modalities including imaging and clinics 86 MCIc 237 MCInc Tabular features $\sim 8\%$ missing features Mean imputation and imputation by the EM Yes $0.670$ n.d. n.d. No [37] sMRI, FDG-PET 76 MCIc 128 MCInc GM volume PET volumes $50\%$ missing PET 3D encoder-decoder network to generate the full PET volume Yes $0.657$ n.d. n.d. No [39] sMRI, FDG-PET Genetics 157 MCIc 205 MCInc ROI based features for PET and MRI, SNPs $51\%$ missing PET Latent representation learning without missing data imputation Yes $0.743$ n.d. n.d. Yes [38] sMRI, FDG-PET 234 MCIc 342 MCInc sMRI and PET volumes $30\%$ missing PET Task-induced pyramid and attention generative adversarial network (TPA-GAN) Yes $0.753$ $0.708$ n.d. No Proposed framework sMRI, fMRI, Genetics 289 MCIc 646 MCInc GM volume sFNC matrix, SNPs $\sim 76\%$ missing fMRI $\sim 45\%$ missing SNPs Multiple cGANs to generate missing latent representation No $0.711\pm 0.01$ $0.610\pm 0.03$ $0.558\pm 0.03$ Yes
CN vs PAT (AD+MCI) classification relying only on sMRI and SNPs, Three classes classification, CN vs MCI vs AD relying on the three modalities
@@ 5.1 Classification performance and missing data management
Focusing on the Task 1, AD detection, our model reached the state-of-the-art performance, obtaining an average testing accuracy of $0.926\pm 0.02$, with the best model reaching an accuracy score of $0.964$. This achievement is even more noteworthy considering that only $12\%$ of the subjects had complete data across all three modalities. Of note, our dataset was constructed considering the subjects that at least had the sMRI modality, however, our proposed framework showcases the capability to potentially impute all three modalities by transferring knowledge across domains, not being limited to having at least one acquisition for all the subjects. Indeed, the injection of the pre-trained generators, that is the two cGANs, allows both to impute the missing rs-fMRI or SNPs from the sMRI and the vice-versa, meaning that it would allow deriving the missing sMRI from the other two modalities if needed. This allows relaxing the requirement of having at least one acquisition for all the subjects which is one of the strengths of the proposed model compared with the state-of-the-art.
Table 3(a) shows the performance of the proposed model compared with the state-of-the-art in detecting AD, considering works allowing multimodality and missing data management. The simplest model [33] proposed a DL-based multimodal method based on sMRI, clinics, and genetics, excluding features if missing in more than $70\%$ of subjects and filling the remaining missing data with zeros. Their classification task was different from ours since they included the MCI subjects in a three-class classification task. While they achieved an accuracy of $0.780$ for the classification between CN, MCI, and AD, the accuracy dropped to $0.630$ when classifying CN from the full patient cohort using only sMRI and genetics. Moreover, the proposed technique has some drawbacks, mainly linked to the biases introduced in the network, as already discussed.
Alternatively, methods to exploit the information of all the available subjects, without explicitly generating the missing modalities were proposed relying on latent representation learning. [35] developed an Auto-Encoder-based multi-view missing data completion framework using sMRI and FDG-PET ROI-based features. Their method achieved a classification accuracy of $0.836$ for classifying CN versus AD, even with $50\%$ missing PET. Similarly, [64] proposed a GAN with an attention layer to generate missing FDG-PET from available MRI features, resulting in a classification accuracy of $0.914\pm 0.19$. While these approaches demonstrated excellent results, their framework was limited in handling arbitrary missing modalities, allowing only the completion of missing PET from available sMRI and FDG-PET features.
Finally, [38] presented a task-induced pyramid and attention generative adversarial network (TPA-GAN) to generate missing FDG-PET data from sMRI, achieving an accuracy of $0.927$. However, the model was trained and tested on two independent databases with different acquisition protocols, which on one side allowed to evaluate the generalizability of the model, but on the other could bias the results due to the different data source. Moreover, it still lacked the ability to reconstruct sMRI from FDG-PET data. Despite the aforementioned limitations, they are currently the state-of-the-art performance for multimodal MCI conversion prediction, accounting for the 30% of missing PET data.
@@
To address classification Task 2, we directly tested the model trained for Task 1 on the MCI cohort. Only the $11\%$ of the dataset shared all the modalities, making it a challenging scenario. Nevertheless, our framework achieved an average accuracy of $0.711\pm 0.01$ for the independent test sets. Although it did not outperform methods specifically trained on this particular task, it demonstrated competitive results. Other approaches have focused on different input data and missing data management to solve this task. For instance, [34] proposed simple machine learning methods with tabular features and limited imputation techniques, achieving an accuracy of $0.670$ to stratify MCI subjects. [37] used a GAN to impute missing PET images from sMRI scans, achieving an accuracy of $0.657$ for discriminating MCInc vs MCIc patients. Furthermore, [39] proposed a framework for projecting original features into a latent representation, resulting in an accuracy of $0.743$ considering a 51% of missing PET. Finally, [38] applied the approach already discussed also to address the classification Task 2, MCI conversion prediction reaching the best accuracy score of $0.753$.
@@
Despite the promising results obtained by generative models for missing data imputation, it is essential to acknowledge their shared limitation: they all relied on having sMRI as a prerequisite to impute PET data, and very few included genetics or rs-fMRI information in their analyses, which have instead been demonstrated as relevant biomarkers for AD [10, 15]. In contrast, our framework does not necessitate a common modality shared by all subjects. Utilizing two cGANs during the generation phase enables to produce the missing latent rs-fMRI and/or SNPs from sMRI. In addition, this process is applicable bidirectionally, allowing the generation of sMRI latent features from either SNPs or rs-fMRI as well. The versatility of this approach opens the way to its generalization to additional modalities through training distinct cGANs and integrating the resulting generators into the complete classification framework. Furthermore, our model’s performance confirms that generative models allow to obtain realistic data and learn nonlinear mappings across the different acquisitions, achieving competitive prediction accuracy also with a substantial proportion of missing modalities.

### v3

@@ 4.3 Group-based statistical IG analysis
Neuroimaging modalities
Figure 6: Boxplots showing the CN, AD, MCInc, and MCIc correctly classified subjects distribution in A. the most relevant cortical and subcortical brain regions resulted from the sMRI-IG, and B. the most relevant pairs of ICs for both intra- and inter-network connections.
@@
Figure 7: The top $15$ biological processes from enrichment analysis reported for the AD, MCIc, and MCInc groups (columns). The colorbar represents the p-value, dark violet the lowest, while the number of genes involved is reported as the bar height. * indicates biological processes surviving Bonferroni correction.
@@
Figure 6A shows the boxplots representing the distribution of the average sMRI-IG values in the most relevant cortical and subcortical brain regions (ROIs), and only those regions where at least the $99\%$ of the subjects had an IG value were retained. The complete acronym list is reported in supplementary Table 2. A general agreement between the CN and the MCInc subjects, as well as between the AD and the MCIc patients is evident in all the considered ROIs, with the AD/MCIc subjects showing a generally higher variance. As expected from the qualitative analysis, the Hipp and the Amy resulted as the highest relevant ROIs, as well as the ones showing the highest group distance, with high positive attributions for the CN and the MCInc, and strong negative attributions for the AD and MCIc. Among the other subcortical ROIs, the Acc showed a notable relevance for all the groups, with the MCInc and CN having negative scores and the MCIc and AD positive ones. Cau and Put had an associated high positive relevance score for AD and MCIc, while almost null scores for the other two groups. Interestingly, among the cortical regions, the TOF was the most relevant for both the MCIc and AD (positive IG scores) and the MCInc and CN (negative IG scores). The largest distance between the group IG score distributions was found for PhGa, followed by the TFCp with the AD and MCIc having high negative values. On the contrary, the OFG, the POpC, and the PaGp showed high positive relevance for AD and MCIc patients associated with negative scores for the CN and MCInc. The Kruskal-Wallis test was performed in order to establish whether differences, separately for each ROI, were significant considering the four groups together. The significant ROIs (all except the Ins and the TP) were further analyzed to investigate the group-related differences through the Wilcoxon Rank Sum test. The most significant differences were found between MCIc and MCInc in the subcortical ROIs, with the lowest p-values being recorded for Hipp and Amy. Moreover, significant differences were also recorded for the temporal (PcC, PhGa, PaGp, TFCp) and occipital (TOF, OFG, POpC) cortical regions, with the most significant difference being recorded for PcC and PhGa. A coherent significance pattern was found between the contrasts MCIc-CN, MCInc-AD, and the CN-AD with generally higher p-values. Of interest, no statistically significant differences were found in the sMRI-IG score distributions between AD and MCIc patients, except for PcC, TFCp, and POpC, while some significant differences were found between CN and MCInc but with relatively high p-values. Supplementary Figure 1A shows the p-values of statistical analysis between groups for each region.
@@
Figure 6B shows the boxplots of fMRI-IG values for the intra- and inter-connections of the four groups. The same pattern of sMRI regarding the similarity between CN and MCInc, and AD and MCIc has been also highlighted for the rs-fMRI intra- and inter-network connections. Among the intra-network connections, the ones related to the sensorimotor area were the most important, such as IC9-IC10, IC9-IC11, IC11-IC12, and IC12-IC13. Regarding the inter-network connections, they showed generally lower relevance with respect to the intra-network ones. The same statistical analysis was carried out for fMRI-IG, considering all the most relevant connections resulting from the fMRI-IG scores for all the considered classes. The connections between IC18181818 (VI) – IC14141414 (SM) and IC44444444 (DM) – IC17171717 (VI) did not reach significance at the Kruskal-Wallis test, therefore they were excluded from the post-hoc analysis. Differently from the sMRI-IG significance pattern, the most significant differences were recorded between the CN-AD and MCInc-AD contrasts, which showed significant differences for almost all the considered connections. In particular, for the MCInc-AD contrast, significant intra-network connections were part of the SM, involving connections in the parietal lobe, and in the VI RSNs. Concerning the inter-network connection, significance was found between ICs in the DM, part of the cingulate gyrus (IC47, IC45), and ICs in the CC or VI RSNs, as well as between ICs in the SM, located in the parietal lobe and ICs in the CC or CB RSNs. MCIc-CN and MCInc-MCIc showed the same significant pattern as the CN-AD contrast, showing generally higher p-values. Of interest, as for the sMRI-IG statistical analysis, no significant differences were detected for the contrasts MCIc-AD and MCInc-CN, except for two connections between CB and SM (IC51-IC9 and the IC51-IC12 respectively for the MCInc-CN and the MCIc-AD). Supplementary Table 3 lists the ICs and corresponding brain regions. The Bonferroni-corrected pairwise tests (Wilcoxon rank sum) are reported in supplementary Figure 1B, separately for intra- and inter-network connections, with the p-values being represented in logarithmic scale.
Genetic modality
@@ 5 Discussion
In this work, we proposed a multimodal generative and interpretable method, which holds the potential of addressing missing data management while focusing on model interpretability. We specifically applied this method to the crucial tasks of segregating AD patients from CN, and assessed its viability in detecting MCI conversion to AD by integrating neuroimaging (sMRI and rs-fMRI) and genetics (SNPs) data. Incomplete data is handled by generating missing modalities in the latent space, obtained after feature reduction. This ensures high accuracy in reconstructing latent features while minimizing computational demands. However, the proposed approach is generalizable to any missing modality, providing a viable and effective solution for overcoming the missing data bottleneck. Furthermore, we conducted a post-hoc interpretability analysis, supplemented by a robust validation step, consolidating the findings’ impact in the field.
@@ 5.1 Classification performance and missing data management
Focusing on Task 1, AD detection, our model reached competitive performance in the state-of-the-art, obtaining an average accuracy of $0.926\pm 0.02$ on the testing set, with the best model reaching an accuracy score of $0.964$. Noteworthy, this result was achieved on a cohort where only the $6.5\%$ of the subjects had complete data across all three modalities. The proposed framework in the same Task 1 also outperformed the two baseline methods with the two imputation strategies, where they achieved an average accuracy of $0.699\pm 0.01$ and $0.663\pm 0.01$ for the ensemble SVM and ensemble random forest with mean imputation strategy, respectively, while accuracies of $0.689\pm 0.02$ and $0.663\pm 0.01$ were obtained using the zero imputation strategy for SVM and random forest, respectively. Our dataset was constructed considering the subjects having at least the sMRI modality, however, the proposed framework showcases the capability to potentially impute all three modalities by transferring knowledge across domains, not being limited to having at least one acquisition for all the subjects. Indeed, the injection of the pre-trained generators, that is the two cGANs, allows both to impute the missing rs-fMRI or SNPs from the sMRI and the vice-versa, meaning that it would allow deriving the missing sMRI from the other two modalities, if needed. This allows relaxing the requirement of having at least one acquisition for all the subjects which is one of the strengths of the proposed model compared with the state-of-the-art.
The simplest model, Venugopalan et al. [38] proposed a DL-based multimodal method based on sMRI, clinical, and genetics data, excluding features if missing in more than $70\%$ of subjects, and filling the remaining missing data with zeros. Their classification task was different from ours since they included the MCI subjects in a three-class classification problem. While they achieved an accuracy of $0.780$ for the classification between CN, MCI, and AD, the accuracy dropped to $0.630$ when classifying CN from the full patient cohort using only sMRI and genetics. Moreover, the proposed technique had some drawbacks, mainly linked to the biases introduced in the network, as already discussed. Alternatively, methods to exploit the information of all the available subjects, without explicitly generating the missing modalities, were proposed relying on latent representation learning. Liu et al. [40] developed an Auto-Encoder-based multi-view missing data completion framework using sMRI and FDG-PET ROI-based features. Their method achieved a classification accuracy of $0.836$ for classifying CN versus AD, even with $50\%$ missing PET. Similarly, Ye and colleagues [45] proposed a GAN with an attention layer to generate missing FDG-PET from available MRI features, resulting in a classification accuracy of $0.914\pm 0.19$. In the same way, Zhang et al. [49] employed a GAN model based on a pyramidal attention mechanism for imputing PET data from sMRI. Then, the generated PET images were fused with the sMRI at pixel-level, and then the same generator used for PET synthesis was employed for feature extraction, extracting a latent representation that fed a dense layer, reaching an accuracy of $0.934$. Zhang et al. [43] developed a GAN-based model, named BPGAN, for generating FDG-PET data from sMRI. They created a 3D multiple convolution U-Net (MCU) generator for PET images, and then they fused the sMRI and the synthetic PET volumes concatenating the data at the pixel-level. Finally, they employed a ResNet-18 for classification achieving an accuracy of $0.981$ in discriminating AD from CN. Gao et al. [42] proposed a multi-level guided GAN (MLG-GAN) and a multimodal transformer (Mul-T) for synthesizing PET and T2-MRI images from sMRI and extracting features for classification from the three modalities, respectively. The MLG-GAN module generated data using voxel-, feature-, and task-level information for a more accurate generation. The Mul-T achieved competitive performance by integrating local and global information from all modalities reaching an accuracy of $0.944$. Tu and colleagues [48] developed a consistent manifold projection GAN (CMPGAN) model for imputing missing PET from sMRI. The CMPGAN integrated a manifold projection operator to map the data in a low dimensional space and a distribution distance measure to ensure that no gradient disappeared during the generator training, hence creating more reliable data. Then, a multilevel multimodal fusion diagnosis network (MMFDN) was developed to diagnose AD by fusing together the synthetic PET and sMRI images. They obtained a final performance in CN versus AD of $0.980\pm 0.003$. Finally, Gao and colleagues [46] presented a task-induced pyramid and attention generative adversarial network (TPA-GAN) to generate missing FDG-PET data from sMRI, achieving an accuracy of $0.927$. Moreover, it still lacked the ability to reconstruct sMRI from FDG-PET data. While these approaches demonstrated excellent results, their framework was limited in handling arbitrary missing modalities.
@@
As mentioned before, the trained model performance was subsequently assessed on Task 2. Only the $7\%$ of the dataset shared all the modalities, making it a challenging scenario. Nevertheless, our framework achieved an average accuracy of $0.711\pm 0.01$ for the independent test sets. Although it did not outperform methods specifically trained on this particular task, it demonstrated competitive results. Other approaches have focused on different input data and missing data management to solve this task. For instance, Ritter and colleagues [39] proposed simple machine learning methods with tabular features and limited imputation techniques, achieving an accuracy of $0.670$ to stratify MCI subjects. Cai et al. [44] used a GAN to impute missing PET images from sMRI scans, achieving an accuracy of $0.657$ for discriminating MCInc vs MCIc patients. Furthermore, Zhou and colleagues [47] proposed a framework for projecting original features into a latent representation, resulting in an accuracy of $0.743$ in case of missing the $51$% of PET data. Gao et al. [46] applied the approach already discussed also to address the classification Task 2, reaching an accuracy score of $0.753$. Gao et al. [42] employed the same framework described in Task 1 of the MCI conversion task. They initially pretrained the models on AD detection, and then they refined the weights on the MCI task, obtaining a final accuracy of $0.778$. Similarly, Tu and colleagues [48] trained and tested their model described for Task 1 in Task 2. They achieved an accuracy of $0.923\pm 0.01$ in the MCI conversion task.
@@
Despite the promising results obtained by the state-of-the-art generative models for missing data imputation, it is essential to acknowledge their shared limitation: they all relied on having sMRI as a prerequisite to impute PET data, and very few included genetics, T2-MRI, or rs-fMRI information in their analyses, which have instead been demonstrated as relevant biomarkers for AD [10, 15]. In contrast, our framework does not necessitate a shared modality across all subjects. The exploitation of two cGANs during the generation phase enables to produce the missing latent rs-fMRI and/or SNPs from sMRI. In addition, this process is applicable bidirectionally, allowing the generation of sMRI latent features from either SNPs or rs-fMRI as well. The versatility of this approach opens the way to its generalization to additional modalities through training distinct cGANs and integrating the resulting generators into the complete classification framework. Furthermore, our model’s performance confirms that generative models allow to obtain realistic data and learn nonlinear mappings across the different acquisitions, achieving competitive prediction accuracy also with a substantial proportion of missing modalities. Supplementary Table 7 shows the performance of the proposed model compared with the state-of-the-art competitors allowing multimodality and missing data management for both Task 1 and Task 2.